A student wishes to determine the molar heat of dissolution of NaOH in water. To do so, the student dissolves 2.00g of NaOH in 100.00g of water that is placed in a coffee cup calorimeter. The initial temperature of the water in the calorimeter was 22.1°C and the final temperature was 26.4°C.
Calculate the molar heat of NaOH dissolution in water in kJ/mol
Answer #1

Heat absorbed by water, q = mC∆T

C is the specific heat capacity of water = 4.186 J/g°C

So, q = 100×4.186×(26.4 - 22.1)

q = 100×4.186×4.3 J = 1799.98 J

Since, 1KJ = 1000J

q = 1799.98/1000 KJ = 1.799 KJ = 1.80 KJ

Now, heat released during dissolution of NaOH is equal to the heat absorbed by water.

i.e. ∆Hdiss = - q (negative sign because heat is released)

∆Hdiss= -1.80 KJ

Molar Mass of NaOH is 40 g/mol

2g NaOH release 1.80 KJ heat

40 g NaOH release 1.80×40/2 KJ = 36 KJ/mol

So, ∆Hdiss = - 36 KJ/mol

Please give feedback !